Compound A decomposes according to the equation

    A(g) ⇌ 2 B(g) + C (g)

A sealed 1.00−L container initially contains 1.84 ×10−3 mol of A(g), 1.29 × 10−3 mol of B(g), and 6.36 ×10−4 mol of C(g) at 100°C.

At equilibrium, [A] is 2.06 × 10−3M.

Find [B] and [C].

Solve for the equilibrium concentrations of B and C.
Answer #1

Ans :

Initial concentration of A = mol / volume (L)

= 1.84 x 10-3 mol / 1.00 L

= 1.84 x 10-3 M

initial concentration of B =  1.29 × 10−3 mol / 1.00 L =  1.29 × 10−3 M

initial concentration of C = 6.36 ×10−4 mol / 1.00 L = 6.36 ×10−4 M

equilibrium concentration of A = 2.06 × 10−3M.

change in concentration of A = 2.06 × 10−3M - 1.84 x 10-3 M = 0.22 x 10-3 M

so change in concentration of B will be = 2 x 0.22 x 10-3 M = 0.44 x 10-3 M and that of C = 0.22 x 10-3 M

now equilibrium concentration of B and C :

[B] = 1.29 × 10−3 M - 0.44 x 10-3 M = 0.85 x 10-3 M

[C] =  6.36 ×10−4 M -  0.22 x 10-3 M = 6.14 x 10-3 M